OpenStudy (anonymous):

8Ix+3/4I <2 A. x<-1 5/6 or 5/6 B .x= -1 or x= 41/3 C. x= -1 5/6 or 4 1/3 D. x= -1 5/6 or X= -1

OpenStudy (xapproachesinfinity):

okay let's play this game hehe the game says first divide both side of the inequality by 8

OpenStudy (anonymous):

is 1.5 the same as 1/2

OpenStudy (xapproachesinfinity):

Now next step is matching this with another rule of the game with is if |x|<a then -a<x<a this is called conjunction

OpenStudy (xapproachesinfinity):

Nope!

OpenStudy (anonymous):

8Ix+3/4I <2 Ix+3/4I <1/4 So x+3/4<1/4 and x+3/4 >-1/4 So x<-0.5 and x>-1 x (-1, -0.5)

OpenStudy (xapproachesinfinity):

1.5 is 3/2 1/2=0.5 surely 3/2>1/2

OpenStudy (xapproachesinfinity):

@WilliamZhang pls do not give complete answers!

OpenStudy (anonymous):

okay so where are we at in the equation

OpenStudy (xapproachesinfinity):

so can you apply the rules of this game to this?

OpenStudy (anonymous):

no...

OpenStudy (xapproachesinfinity):

we said first divide by 8 both sides then look at what case do we have

OpenStudy (anonymous):

4

OpenStudy (xapproachesinfinity):

if |x|<a conjunction then -a<x<a

OpenStudy (xapproachesinfinity):

a must be positive though to apply this! just a warning

OpenStudy (anonymous):

so then it would be C

OpenStudy (xapproachesinfinity):

man! do the word and see what you get

OpenStudy (xapproachesinfinity):

work*

OpenStudy (xapproachesinfinity):

Do you want to play the game or not? if not! no answer for you

OpenStudy (xapproachesinfinity):

think of this as a game will ya!

OpenStudy (anonymous):

well im trying to figure out but I don't think the game is working

OpenStudy (xapproachesinfinity):

well the game is good to me! we started at \(8|x+\frac{3}{4}|<2\) dividig by 8 \( \large\frac{8}{8}|x+\frac{3}{4}|<\frac{2}{8}\) 8/8 is just 1 so \( |x+\frac{3}{4}|<\frac{1}{4}\) 2/8 is 1/4 the game said so haha no you have conjuction

OpenStudy (anonymous):

Simplify the inequality , Then,if |A|<B then -B<A<B. Do the simple algebra. Thats it.

OpenStudy (xapproachesinfinity):

now* not "n" sorry

OpenStudy (xapproachesinfinity):

well! how about you continue this game?

OpenStudy (anonymous):

k

OpenStudy (xapproachesinfinity):

so according to our game rule what should we do next?

OpenStudy (anonymous):

add

OpenStudy (xapproachesinfinity):

add what i will repost the games rule if |x|<a, a is positive then -a<x<a

OpenStudy (xapproachesinfinity):

I keep saying this many times you pay attention to it^_^

OpenStudy (xapproachesinfinity):

a is our case is 1/4 x is our case is x+3/4

OpenStudy (xapproachesinfinity):

can you see it now?

OpenStudy (anonymous):

so divide the 3/4 to 1/4

OpenStudy (xapproachesinfinity):

in our case* sorry for bad miss spelling

OpenStudy (xapproachesinfinity):

No! look |x|<a, so is becomes -a<x<a

OpenStudy (xapproachesinfinity):

instead of a put 1/4 instead of x put x+3/4

OpenStudy (anonymous):

so x+ 3/4 <1/4

OpenStudy (xapproachesinfinity):

but i want you to remember that a should be positive to do this rule otherwise we would have done a different thing! just keep this in mind for future problems^_^

OpenStudy (xapproachesinfinity):

what about the other side of the inequality? -a<x<a look and see if you did the right thing?

OpenStudy (anonymous):

-1< x<-0.5 , A. x<-1 5/6 or 5/6 B .x= -1 or x= 41/3 C. x= -1 5/6 or 4 1/3 D. x= -1 5/6 or X= -1 so no above is correct.

OpenStudy (xapproachesinfinity):

@WilliamZhang pls no answers don't confuse our friend we trying to learn this. we are not trying to just solve it?

OpenStudy (xapproachesinfinity):

we teach me how i fish not give me the fish lol if you give me the fish next time you are not around i cannot get the fish

OpenStudy (anonymous):

Sorry, I am just confused about the options.

OpenStudy (xapproachesinfinity):

we will see it through if there is anything but let's keep it simple for now

OpenStudy (anonymous):

@WilliamZhang A <-1 or X >-1/2 B -5/32<x<5/32 C -1 ,x, -1/2 D is <-5/32 or x> 5/32

OpenStudy (anonymous):

sorry I typed out something different

OpenStudy (xapproachesinfinity):

i figured out that! So i was gonna ask you to relook at but that was not important for now!

OpenStudy (xapproachesinfinity):

once we get the right thing you would look at the answer and you will say look that is the what i got too in this game lol

OpenStudy (anonymous):

so its positive we know that

OpenStudy (xapproachesinfinity):

now pls go back to where we are you said x+3/4<1/4 but i said that's not complete?

OpenStudy (anonymous):

so it would be B

OpenStudy (anonymous):

8|A|<2 |A|<2/8 |A|<(2*1)/(2*4) |A|<1/4 U get it? Then, if |A|<B then -B<A<B. U have to remember this.

OpenStudy (xapproachesinfinity):

Again you said be! i don't really care whether is A or B or C or D our main focus is to solve this.

OpenStudy (xapproachesinfinity):

and learn how to do it!

OpenStudy (anonymous):

1/4< x=3/4 <14

OpenStudy (xapproachesinfinity):

in fact i haven't looked at the solution yet haha.

OpenStudy (anonymous):

so we would divde

OpenStudy (xapproachesinfinity):

that is incorrect! relook at it again?

OpenStudy (anonymous):

that's how I just learned it

OpenStudy (xapproachesinfinity):

i said -a<x<a

OpenStudy (anonymous):

-1/4 <3/4x <1/4

OpenStudy (xapproachesinfinity):

that is okay! learning take time! don't be afraid! ashamed or anything! okay just continue learning. no matter say i have to learn it

OpenStudy (xapproachesinfinity):

you have x+3/4 how did it become 3/4x?

OpenStudy (anonymous):

that's what I meant

OpenStudy (xapproachesinfinity):

the other parts are fine^_^

OpenStudy (xapproachesinfinity):

okay! be very precise! math is precise^_^ so -1/4<x+3/4<1/4

OpenStudy (anonymous):

so divide -1/4and 3/4

OpenStudy (xapproachesinfinity):

now we want to separate x by it self so we need to subtract 3/4 but we need to subtract it from all side

OpenStudy (anonymous):

so subtract 3/4 by 1/4

OpenStudy (xapproachesinfinity):

Not divide! we don't have multiplication to divide we have addition here so subtraction is needed

OpenStudy (xapproachesinfinity):

look at this -1/4-3/4<x+3/4-3/4<1/4-3/4

OpenStudy (xapproachesinfinity):